Sunny Sunday – University of Copenhagen

24 May 2009

Sunny Sunday

Setting up flags on new skiway.

Setting up flags on new skiway.

This Sunday morning, the main power of camp was disconnected, as Bruno exchanged the main power distribution panel on the generator hut. It took from 7 AM to about 1 PM. In the mean time, our 16 kW backup generator powered the main dome. Everything went fine, and as soon as the exchange was completed, we were able to get back to work in the trenches. It was a fine day again, and work progressed at a steady pace. In the last days, the amount of snow blocks that has been taken from excavations in the science trench, has grown to a formidable hill. Several camp members used the nice, sunny evening to build a snow hut (to call it an igloo would be an insult to igloos). They were having great fun.

What we have done today:

  1. Drilling.
  2. Exchanging power distribution outlet on generator hut.
  3. Excavating a cave for physical properties cabin.
  4. Measuring ECM on new core from this year. A match of two volcanic layers between the 2008 core and the 2009 core has been established.
  5. Working on roof for elevator and staircase.
  6. Removing large snow drifts in camp.
  7. Setting up tables and shelves in CFA lab.
  8. Taking down flags on ½ of the old skiway and setting them up on the new skiway.
  9. Working on assembling the new NEEM drill.

Ad.1: This is how the drillers' report looks today:
“Activities today took place in three parallel tracks. Trevor, Krissy, and Nobby continued with the EPICA drill for five runs using the same configuration each time (4.5 mm pitch shoes, 134 mm diameter cutter configuration, antitorque skates set to 38 mm pilehøjde). Smooth drilling was the norm with one antitorque slip when motor current limit was reached during the last run. At present we limit the length of the runs to two metes to accommodate core troughs for storage while the loggers establish the new depth reference based on ECM matching. We tested the cable cleaner device designed by Jakob and early indications are that it effectively dries the cable of excess liquid. One great property of the drilling liquid is that excess chips we do not collect float and therefore do not impair drilling down hole. When necessary, these chips are easily collected with the bottom filter. Meanwhile Steffen and Sigfus continue to put together the new long drill. Finally, Jakob continues to repair the winch control inverters and has remounted the original inverters into their original boxes.
Driller’s depth is now 134.44 meters.”

Ad.4: We have matched the ECM signal from last year’s pilot hole to the new main hole. As the drill diverted away from the old hole, we now have an overlap of 5 meter between the two cores. A match of two volcanic signals (1646 AD and 1641 AD) in both cores, has resulted in a depth scale adjustment from drillers depth to core depth of 1 cm!! The new NEEM main core begins with Bag 180 (98.45 m below 2008 surface).

Weather: Clear all day, -22°C to -11°C, 5-6 knots mainly from S and SSE.
Visibility: Unrestricted. Beautiful.

FL, J.P. Steffensen

← Previous entry   Next entry →